OpenStudy (anonymous):

What is ms^-1 in terms of velocity? It's a unit right, but whats the -1 for?

OpenStudy (anonymous):

\[Example: 2ms ^{-2}\]

OpenStudy (anonymous):

sorry... I meant to write -1

OpenStudy (anonymous):

In fact formula of velocity is \[v = \frac{ d }{ t }\]

OpenStudy (anonymous):

And "d' stands for displacement or distance... SI unit of d is meter... and " t " stands for time.. SI unit of time is Second...

OpenStudy (anonymous):

yeah but like the velocity can be written km/h or m/s but why ms^-1

OpenStudy (anonymous):

velocity in terms of unit is.. \[v = \frac{ meter }{ second }\]

OpenStudy (anonymous):

but why the -1?

OpenStudy (anonymous):

So unit of velocity is \[\frac{ m }{ s }\]

OpenStudy (anonymous):

oh!!! haha!! lightbulb moment

OpenStudy (anonymous):

\[ms ^{-1}\]

OpenStudy (vincent-lyon.fr):

you can also write \(km.h^{-1}\), the negative power is not restricted to \(m.s^{-1}\)

OpenStudy (anonymous):

got it. Index laws

OpenStudy (anonymous):

thanks!!!

mathslover (mathslover):

\(\textbf{anything}^{-1} = \cfrac{1}{\textbf{anything}}\)

OpenStudy (anonymous):

can't believe I forgot that!!

OpenStudy (anonymous):

You're welcome.